# Sensing with near-infrared laser trapped fluorescent nanodiamonds

### v1

@@ III.1 Photoluminescence Spectrum Analysis
We conducted PL spectroscopy of FNDs with varying NIR power. The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium.
Effect of NIR Laser Power on PL Spectrum Intensity
We exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW and for each power observe the PL spectrum from 550 nm to 750 nm with our spectrometer.
Figure 4: Effect of CW NIR laser on dry FND emission spectrum. Optical emission spectrum of dry FND on glass under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
Our PL measurements revealed significant alterations in the spectral intensities of both zero-phonon lines (ZPL) from the two NV charge states upon CW NIR laser exposure. Notably, as a function of NIR laser power there was a slightly more pronounced reduction in the spectral intensity of NV- compared to NV0. As shown in Fig. 4, the ZPL intensity of NV0 at 575 nm decreased by 29.5% and 32.4%, while the NV- ZPL intensity at 637 nm decreased by 52.8% and 62.3% when the CW NIR laser power was increased from 0 mW to 30 mW and 60 mW, for the condition with FNDs dry on glass. For the fluorescence spectrum (like Fig. 4), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water, cf. Fig. 5. Quantitatively, we observe a 1.2% general intensity difference from 30 mW to 60 mW for FNDs on glass with water while the same figure is 0.7% for FNDs trapped in water.
Figure 5: Effect of CW NIR laser on trapped FND emission spectrum. Optical emission spectrum of FND trapped in water CW NIR laser at 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
To determine the ratio of NV0 to NV- emission intensities under CW NIR laser powers of 0 mW, 30 mW, and 60 mW we integrated the emission intensities over the respective spectral ranges for NV0 (550–620 nm) and NV- (655–750 nm), but compared to the procedure of Ref. [13], we applied a normalization factor to account for the difference in spectral ranges:
$\frac{\text{NV}^{0}}{\text{NV}^{-}}=\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{750}I_{\lambda}}\right)\times\frac{95}{70}$ (1)
We observe that the NV0/NV- emission ratio stays the same for trapped FNDs, 0.45 for both 30mW and 60mW NIR laser power. In contrast, it rises from 0.39 without NIR irradiation to 0.51 with 60 mW NIR laser power for dry FND on glass. All results are displayed in Fig. 10.
@@ III.2 Optically Detected Magnetic Resonance Measurements
We further examined the influence of CW NIR laser power on the ODMR resonance frequency and contrast for FNDs dried on glass substrates and those trapped in water.
Effect of CW NIR Laser Power on the Zero Field Splitting of the NV Center
We exposed the FNDs to CW NIR laser powers of 0 mW, 30 mW, and 60 mW, and measured the resulting ODMR spectra. Our primary observation was a shift in the resonance frequency with increasing CW NIR laser power, as illustrated with an example in Fig. 6.
@@
Figure 6: Effect of CW NIR laser power on ODMR spectrum of dry FND. ODMR spectra of dry FNDs under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power. The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed.
@@
For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies. Specifically, the resonance frequency shift increased in absolute value from $-0.31\pm 0.54$ MHz at 0 mW to $-5.83\pm 046$ MHz at 60 mW of CW NIR power; both values given relative to the expected 2.87 GHz. In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $-0.84\pm 0.21$ MHz at 30 mW and $-1.58\pm 0.33$ MHz at 60 mW (relative to the expected 2.87 GHz). Examples are shown in Fig. 6 and Fig. 7 while combined results are displayed in Fig. 11. As before, for the similar experiment when water was added to the same FND that remained stuck on the glass slide, the ODMR spectra show a trend very similar to Fig. 6 although the difference between 30 mW and 60 mW of NIR power is less pronounced (data not shown).
Figure 7: Effect of CW NIR laser power on ODMR spectrum of trapped FND. ODMR spectra of FNDs trapped in water under CW NIR laser exposure at 30 mW (orange) and 60 mW (red) of 1064 nm laser power. The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed.
III.3 T1 Relaxometry Measurements
@@
We conducted $T_{1}$ relaxometry experiments under different NIR laser power. As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the T1 relaxometry measurements with NIR irradiation. The samples were exposed to NIR laser powers of 0 mW, 30 mW, and 60 mW, and the corresponding $T_{1}$ values were measured. For these experiments, the pulsing sequence in Fig. 3 was applied.
Effect of NIR Laser Power on $T_{1}$ Relaxation Times
First, we investigate how the T1 value of dry FNDs is affected by NIR irradiation. For FNDs dried on glass substrates, we observed a significant decrease in the $T_{1}$ relaxation times with increasing NIR laser power. As shown in the example in Fig 8, the mean $T_{1}$ value decreased from $241\pm 21~{}\mu$s at 0 mW NIR power to $136\pm 19~{}\mu$s at 60 mW NIR power. When water was added to the same FND that remained stuck on the glass slide, the $T_{1}$ values showed very similar changes. The mean $T_{1}$ value decreased from $237\pm 25~{}\mu$s at 0 mW NIR power to $154\pm 17~{}\mu$s at 30 mW NIR power and $143\pm 21~{}\mu$s at 60 mW NIR power. (data not shown).
@@
Figure 8: Effect of NIR laser power on $T_{1}$ relaxation times of dry FNDs. $T_{1}$ relaxometry measurements for FNDs dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength. Pulsing as in Fig 3.
For FNDs trapped in water, the decrease in $T_{1}$ due to NIR irradiation was significantly less. As depicted in the example in Fig. 9, we do not observe any significant change in the 2 curves. The $T_{1}$ values for trapped FNDs were $179\pm 8~{}\mu$s at 30 mW and $176\pm 8~{}\mu$s at 60 mW, showing a negligible decrease compared to the values at 0 mW, with all results summarized in Fig. 11.
Figure 9: Effect of NIR laser power on $T_{1}$ relaxation times of trapped FNDs. $T_{1}$ relaxometry measurements for FNDs trapped in water under pulsed NIR laser powers of 30 mW (orange) and 60 mW (red) at 1064 nm wavelength.
III.4 Combined Statistical Analysis
In this section, we present a comprehensive statistical analysis of our experimental results.
Relationship Between NV0/NV- Ratio and ODMR Frequency Shift
Figures 10.a, 10.c, 11.a and 11.b illustrate the dependence of the NV0/NV- fluorescence intensity ratio and the ODMR resonance frequency shift for FNDs under varying CW NIR laser powers.
For FNDs dried on glass substrates, we observe a significant increase in the NV0/NV- fluorescence intensity ratio and a corresponding shift in the ODMR resonance frequency with increasing CW NIR power from 0 mW to 60 mW.
In contrast, for FNDs trapped in water, the NV0/NV- fluorescence intensity ratio and ODMR frequency shift exhibit much smaller changes with increasing CW NIR power. This suggests that the aqueous environment mitigates the effects of CW NIR irradiation.
Correlation Between $T_{1}$ Relaxation Times and ODMR Frequency Shift
Figures 10.b, 10.d, 11.a and 11.c show the relationship between the $T_{1}$ relaxation times and the ODMR resonance frequency shift for FNDs under varying NIR laser powers. $T_{1}$ was obtained with the pulse sequence as in Fig 3 while the ODMR was obtained with CW NIR irradiation.
For FNDs dried on glass substrates, there is an effect between $T_{1}$ times and ODMR frequency shifts with increasing NIR power. The significant decrease in $T_{1}$ times and increase in frequency shift suggest that NIR irradiation enhances spin relaxation processes.
For FNDs trapped in water, the $T_{1}$ relaxation times remain relatively constant, and the ODMR frequency shifts exhibit minimal changes with increasing NIR power.
@@ III.4 Combined Statistical Analysis
Overall, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts. Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf. Fig. 3 and Ref. [19], potential detrimental effects of NIR on $T_{1}$ relaxation can be significantly mitigated, even with FNDs dry on glass.
@@
Figure 10: a) Correlation between NIR laser power, ODMR resonance frequency shift and NV0/NV- fluorescence intensity ratio for FNDs trapped inside water. b) Correlation between NIR laser power, ODMR resonance frequency shift and T1-relaxometry times for FNDs trapped inside water. c) Correlation between NIR laser power, ODMR resonance frequency shift and NV0/NV- fluorescence intensity ratio for FNDs dry on glass. d) Correlation between NIR laser power, ODMR resonance frequency shift and T1-relaxometry times for FNDs dry on glass. Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation. CW NIR for ODMR and spectra, pulsed NIR as in Fig 3 for T1-relaxometry. In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and no such trend for FNDs in the optical trap in water.
@@
Figure 11: Summary of experimental results, showing variation for individual nanodiamonds. (a) ODMR resonance frequency shift, (b) NV0/NV- fluorescence intensity ratio, and (c) $T_{1}$ relaxometry times as functions of NIR laser power. Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
Statistical Analysis Using ANOVA
@@
To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{1}$ relaxation times, ODMR frequency shifts, and NV0/NV- fluorescence intensity ratios. The ANOVA results are summarized in Table 1.
@@
Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water. The p-values indicate the significance of differences observed across NIR power levels for $T_{1}$ relaxation times, ODMR frequency shifts, and NV0/NV- ratios.
@@
$T_{1}$ Relaxation Time $<0.0001$
ODMR Frequency Shift $<0.0001$
NV0/NV- Ratio 0.0024
@@
NV0/NV- Ratio 0.371
For FNDs dried on glass substrates, the very low p-values (p $<0.01$) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant. This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
For FNDs on glass substrates after adding water, the medium p-values ($0.20.01<p<$ 0.2) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant. In fact, the difference between no NIR laser and NIR laser is significant but the differences between 30 mW and 60 mW power of NIR laser are not significant.
In contrast, for FNDs trapped in water, the p-values are high (p $>0.07$), suggesting that the differences observed across different NIR power levels are not statistically significant. This reinforces the conclusion that the aqueous environment mitigates the effects of NIR irradiation on the NV centers.
III.5 Charge State Dynamics
To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [32]. This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation.
The NV center can exist in two charge states: the negatively charged NV- and the neutrally charged NV0. Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable singlet states. The model considers 8 states as illustrated in the schematic energy level diagram in Fig. 12.
The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [32]. The rates of change had to be adapted to our experimentation with green 532 nm and NIR 1064 nm laser.
@@ III.5 Charge State Dynamics
Figure 12: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), singlet states (P5 and P6), and the transitions between them. Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
Population Dynamics of NV- and NV0 Centers
@@
However, as the CW NIR power increases beyond this threshold, the trend reverses. The NV- population begins to decrease, and the NV0 population increases. At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly. This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0. The threshold depends on the recombination rate, ionization rate and on the initial population of each state.
Fluorescence Intensity and NV0/NV- Ratio
@@
The ratio of NV0 to NV- fluorescence intensities ($F_{\text{NV}^{0}}/F_{\text{NV}^{-}}$) increases with CW NIR power. At low CW NIR powers, this ratio is low due to the predominance of NV- fluorescence. As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission. This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
Consistency with Experimental Observations
The trends observed in the simulation align with the experimental results. The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity and NV0/NV- fluorescence intensity ratio observed experimentally. The simulation provides a quantitative framework to understand how NIR illumination influences the charge state dynamics and, consequently, the optical and spin properties of NV centers.
Figure 13: Total fluorescence intensity from NV centers as a function of NIR laser power. (Black: Theory, Red: Experimental observations.
III.6 Temperature Simulation
To model the temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer. The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air). The key processes considered are:
  1. 1.
    Energy Absorption: The FND absorbs a portion of the incident laser power, leading to an increase in its internal energy and temperature.
  2. 2.
    Heat Loss: Heat is lost from the FND to the surrounding medium through conduction and convection, depending on the thermal properties of the medium.

### v2

@@ III.1 Photoluminescence Spectrum Analysis
As the NV neutral and negative charge states have distinct optical emission spectra, we conducted PL spectroscopy of FNDs at varying NIR power to investigate its impact. The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium. We record the PL spectrum in the range between 550 nm to 745 nm with the spectrometer for each of the different values of the CW NIR laser power.
Figure 3: Effect of CW NIR laser on dry FND emission spectrum. Optical emission spectrum of one FND, dry on glass, under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
Upon CW NIR laser light exposure, our PL measurements revealed a significant drop in the total PL intensity. Notably, as a function of NIR laser power there was a slightly more pronounced reduction in the spectral intensity of NV- compared to NV0. For the fluorescence spectrum (like Fig. 3), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water.
To quantify the changes in relative charge state emission intensities under CW NIR laser irradiation, we determine the ratio $\sigma_{0-}$ between NV0 and NV- emission intensities by integrating over the spectral ranges for NV0 (550–620 nm) and NV- (655–745 nm), respectively. Compared to the analysis used in Ref. [Sow2020], we applied a normalization factor 19/14 to account for the difference in the relative spectral ranges:
$\sigma_{0-}\equiv\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{745}I_{\lambda}}\right)\times\frac{19}{14}$ (1)
For FNDs trapped with NIR powers of 30 mW and 60 mW, we observe an averaged similar emission ratios $\sigma_{0-}$ of $0.48\pm 0.09$. In contrast, for FNDs dried on glass, we measure an emission ratio of $0.43\pm 0.11$ without NIR irradiation that increases to $0.57\pm 0.15$ for a NIR laser power of 60 mW NIR. All results are displayed in Fig. 6 parts a) and c).
@@ III.2 Optically Detected Magnetic Resonance Measurements
In the next step, we elaborate on the influence of CW NIR laser light on the ODMR resonance frequency and contrast, revealing the impact of the NIR trapping laser on the NV electron spin properties.
The ODMR spectra were acquired in continuous-wave (CW) mode with constant intensities of both the 532 nm pump and 1064 nm trapping lasers and the microwave source, while sweeping the frequency of the microwave in steps of 0.3 MHz. Generally and for all FNDs, we observe a redshift in the resonance frequency with increasing CW NIR laser power, as illustrated in Fig. 4 by the spectra from one FND with 0 mW and 60 mW NIR power.
@@
Figure 4: Effect of CW NIR laser power on ODMR spectrum of dry FND. ODMR spectra of one FND under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power. The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed. The power of the 532 nm laser was 0.7 mW, measured before the objective lens, and the microwave amplifier was set to -2dB.
@@
For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies. Taking 2.87 GHz as reference value for the zero field splitting, the resonance frequency shift increased in absolute value from $0.31\pm 0.54$ MHz at 0 mW to $5.83\pm 046$ MHz at 60 mW of CW NIR power. In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $0.84\pm 0.21$ MHz at 30 mW and $1.58\pm 0.33$ MHz at 60 mW. Examples are shown in Fig. 4 and combined results displayed in Fig. 7. As before, for the similar experiment when water was added to the same FND that remained stuck on the glass slide, the ODMR spectra show a trend very similar to Fig. 4 although the difference between 30 mW and 60 mW of NIR power is less pronounced (data not shown).
[…]
III.3 Relaxometry Measurements
@@
Last, we conducted relaxometry experiments under different NIR laser power, to extract values for the corresponding $T_{\rm relax}$. As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the relaxometry measurements with NIR irradiation. For these experiments, the all-optical pulsing sequence in Fig. 2b) was applied.
[…]
@@
Figure 5: Effect of NIR laser power on $T_{\rm relax}$ relaxation times of dry FNDs. $T_{\rm relax}$ relaxometry measurements for one FND dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength. Pulsing as in Fig 2b).
[…]
@@
Overall, as showed in Fig. 6 and Fig. 7, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts. Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf. Fig. 2b) and Ref. [Russell2018a], potential detrimental effects of NIR on $T_{\rm relax}$ relaxation can be significantly mitigated.
@@
Figure 6: a) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs trapped inside water. b) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs trapped inside water. c) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs dry on glass. d) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs dry on glass. Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation. CW NIR for ODMR and spectra, pulsed NIR as in Fig 2b) for fluorescence relaxometry. In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and and no or much less pronounced trend for FNDs in the optical trap in water.
@@
Figure 7: Summary of experimental results, showing variation for individual nanodiamonds. (a) ODMR resonance frequency shift, relative to the expected resonance at 2.87 GHz, (b) fluorescence intensity ratio $\sigma_{0-}$, and (c) $T_{\rm relax}$ relaxometry times as functions of NIR laser power. Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
@@
To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios. The ANOVA results are summarized in Table 1 (one-way Kruskal–Wallis test, $\alpha=0.05$).
@@
Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water. The p-values indicate the significance of differences observed across NIR power levels for $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
@@
$T_{\rm relax}$ Relaxation Time <0.0001<0.0001
ODMR Frequency Shift <0.0001<0.0001
$\sigma_{0-}$ Ratio 0.0024
@@
$\sigma_{0-}$ Ratio 0.371
For FNDs dried on glass substrates, the very low p-values (p <0.01<0.01) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant. This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
For FNDs on glass substrates after adding water, the mean p-values ($0.01<p<0.2$) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant. In fact, the difference between no NIR laser and NIR laser is significant but the differences between 30 mW and 60 mW power of NIR laser are not significant.
In contrast, for FNDs trapped in water, the p-values are high (p >0.07>0.07), suggesting that the differences observed across different NIR power levels are not statistically significant. This reinforces the conclusion that the aqueous environment mitigates the effects of NIR irradiation on the NV centers.
III.4 Model of Charge State Dynamics
To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [Qian2022], supplemented with insights from [Meirzada2018]. This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation.
The NV center can exist in two charge states: the negatively charged NV- and the neutrally charged NV0. Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable quartet states. The model considers 8 states as illustrated in the schematic energy level diagram in Fig. 8.
The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [Qian2022]. The rates of change were adapted to our experimental situation with just two lasers, the green 532 nm and the NIR 1064 nm laser, following [Meirzada2018]. Values used in this work for the rates are provided in the appendix, Table 3.
@@ III.4 Model of Charge State Dynamics
Figure 8: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), quartet states (P5 and P6), and the transitions between them. Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
@@
However, as the CW NIR power increases beyond this threshold, the trend reverses. The NV- population begins to decrease, and the NV0 population increases. At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly. This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0.
@@
The $\sigma_{0-}$ ratio increases with CW NIR power. At low CW NIR powers, this ratio is low due to the predominance of NV- fluorescence. As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission. This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
The trends observed in the simulation align with the experimental results. The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity (Fig. 9) and $\sigma_{0-}$ ratio (Fig. 6a,c)) observed experimentally.
Figure 9: Total fluorescence intensity from NV centers as a function of NIR laser power. (Black: Theory, Red: Experimental observations with errorbars, obtained as a statistical average over results from 10 different FNDs with $\sigma_{0-}$ within the range of $\sigma_{0-}=0.42\pm 0.04$.)
III.5 Temperature Simulation
To model the temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer. The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air).
To quantify the temperature rise produced by the 1064 nm laser on a spherical fluorescent nanodiamond (radius $R=35\;\text{nm}$), we solve Fourier’s heat‑conduction equation with an internal absorption source inside the particle and pure conduction in the surrounding infinite medium (water or air):
$\rho_{d}c_{d}\frac{\partial T_{d}}{\partial t}$ $=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\left(k_{d}r^{2}\frac{\partial T_{d}}{\partial r}\right)+q_{\mathrm{abs}}(r),$ $0\leq r<R,$
$\rho_{m}c_{m}\frac{\partial T_{m}}{\partial t}$ $=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\left(k_{m}r^{2}\frac{\partial T_{m}}{\partial r}\right),$ $r>R,$
subject to the interface and far‑field conditions
$T_{d}(R,t)=T_{m}(R,t),$
$k_{d}\!\left.\frac{\partial T_{d}}{\partial r}\right|_{r=R}=k_{m}\!\left.\frac{\partial T_{m}}{\partial r}\right|_{r=R},\qquad\lim_{r\to\infty}T_{m}(r,t)=T_{\infty}.$
Here $\rho$, $c$, and $k$ are the density, specific heat, and thermal conductivity of the nanodiamond and the surrounding medium. The laser absorption is treated as a uniform volumetric heat source inside the FND,
$q_{\mathrm{abs}}(r)=\begin{cases}\dfrac{P_{\mathrm{abs}}}{\tfrac{4}{3}\pi R^{3}},&r<R,\\[6.0pt] 0,&r\geq R,\end{cases}$
where $P_{\mathrm{abs}}=\sigma_{\mathrm{abs}}I_{0}$ is the absorbed optical power for a beam of intensity $I_{0}$ and absorption cross‑section $\sigma_{\mathrm{abs}}$.